# Symmetric Entropy-Constrained Video Coding for Machines

### v1

@@ Abstract
As video transmission increasingly serves machine vision systems (MVS) instead of human vision systems (HVS), video coding for machines (VCM) has become a critical research topic. Existing VCM methods often bind codecs to specific downstream models, requiring retraining or supervised data and thus limiting generalization in multi-task scenarios. Recently, unified VCM frameworks have employed visual backbones (VB) and visual foundation models (VFM) to support multiple video understanding tasks with a single codec. They mainly utilize VB/VFM to maintain semantic consistency or suppress non-semantic information, but seldom explore how to directly link video coding with understanding under VB/VFM guidance. Hence, we propose a Symmetric Entropy-Constrained Video Coding framework for Machines (SEC-VCM). It establishes a symmetric alignment between the video codec and VB, allowing the codec to leverage VB’s representation capabilities to preserve semantics and discard MVS-irrelevant information. Specifically, a bi-directional entropy-constraint (BiEC) mechanism ensures symmetry between the process of video decoding and VB encoding by suppressing conditional entropy. This helps the codec to explicitly handle semantic information beneficial for MVS while squeezing useless information. Furthermore, a semantic–pixel dual-path fusion (SPDF) module injects pixel-level priors into the final reconstruction. Through semantic-pixel fusion, it suppresses artifacts harmful to MVS and improves machine-oriented reconstruction quality. Experimental results show our framework achieves state-of-the-art (SOTA) in rate–task performance, with significant bitrate savings over VTM on video instance segmentation (37.41%), video object segmentation (29.83%), object detection (46.22%), and multiple object tracking (44.94%). We will release our code.
@@ I Introduction
Figure 1: Overview of different methods. $I(\cdot)$ denotes mutual information, $H(\cdot)$ denotes Shannon entropy, and $\Uparrow$” denotes that the optimization goal is to maximize the value. $\mathcal{E}_{\textit{visual}}$ denotes the encoder of visual backbone (VB) or visual foundation model (VFM). (a) Standard or neural video codecs [1, 2] are optimized for HVS, which makes the decoded videos $\hat{x}$ suboptimal for machine vision. (b) Each neural video/feature codec [3, 4] is optimized for a specific machine vision task. (c) The unified codec [5, 6, 7, 8] is optimized with the assistance of the VB or VFM to reconstruct videos suitable for multiple downstream tasks. (d) Our unified compression framework aligns the video coding process with the video understanding process from the entropy perspective and squeezes machine-vision-irrelevant information during the process of video reconstruction.
Digital video plays a crucial role in our daily lives, comprising a significant portion of data storage. To efficiently store and transmit this massive amount of data, video coding techniques have been developed, primarily designed to meet the requirements of the human visual system (HVS). For videos aimed at HVS, such as movies and short clips, it is essential to preserve visual details perceptible to humans during compression. Meanwhile, videos collected for machine vision tasks, such as surveillance analysis [9] and facial recognition [10], typically require only part of the video information, rather than the complete visual details [11, 12]. Currently, neural video coding frameworks have evolved significantly and offer excellent compression performance [13, 2, 14, 15, 16]. However, these methods are optimized primarily for HVS-related metrics and do not fully satisfy the requirements of machine vision systems (MVS) in terms of video understanding (e.g., detection, segmentation, and tracking) [17], as shown in Fig. 1(a). Therefore, it is essential to develop video coding for machines (VCM) methods and bridge the gap between video coding and video understanding.
To reconstruct machine-vision-friendly videos, most existing methods [18, 19, 4] bridge the upstream codec with specific downstream task models and jointly optimize with the loss functions of different tasks, as shown in Fig. 1(b). However, this requires retraining the entire upstream codec or downstream models for each new task, resulting in significant computational and time costs. Meanwhile, some parameter-efficient fine-tuning (PEFT) strategies are introduced to adapt codecs to new downstream tasks with lower cost [3]. Nonetheless, most of them still rely on the labeled data and supervised training process for specific downstream tasks, limiting their generalization in multi-task scenario.
@@
To address the above issues, we propose a Symmetric Entropy-Constrained Video Coding framework for Machine (SEC-VCM). It supports multiple downstream tasks with a single unsupervised training process and further eliminates machine-vision-irrelevant information by leveraging existing visual backbones, as shown in Fig. 1(d). Its key principle is to bridge the gap between the video coding and understanding processes from an entropy perspective, explicitly guiding the video codec to handle only information beneficial for video understanding while discarding irrelevant information. Specifically, an information-alignment-based semantic decoder with a bi-directional entropy-constraint (BiEC) mechanism is proposed to make the video decoding process of the codec symmetric and consistent with the encoding process of the visual backbone. In addition, a semantic-pixel dual-path fusion (SPDF) module is designed to incorporate HVS-related priors into the machine-oriented reconstruction process. It helps reduce artifacts harmful to MVS and fulfill the frame content through a gated fusion mechanism. The experiments on four machine vision tasks demonstrate that the proposed SEC-VCM achieves state-of-the-art (SOTA) rate-task performance. Compared to VTM, our method achieves significant bitrate reductions, with savings of 37.41% for video instance segmentation, 29.83% for video object segmentation, 46.22% for object detection, and 44.94% for multiple object tracking.
@@ II-A Neural Video Coding
Most existing neural video coding methods [27, 28, 29, 14, 30, 31, 32, 15, 16, 33, 34] follow the motion-then-residual paradigm and mainly include inter-prediction and residual/context compression. Lu et al. [27] proposed the first end-to-end video coding framework DVC, which uses the optical flow [35] for inter-prediction. Agustsson et al. [30] proposed the Scale Space Flow and added a Gaussian dimension to create more reference pixels for inter-frame prediction. Hu et al. [36] proposed FVC to convert the video from the pixel domain to the feature domain and used deformable convolution for ME-MC in the feature domain. Li et al. [28] improved the concept of predictive coding and proposed the conditional coding paradigm for better compression performance, termed DCVC. Sheng et al. [37] upgraded the DCVC and proposed DCVC-TCM with a multi-scale conditional coding to better remove temporal redundancy. Furthermore, many methods [15, 2, 38] introduce long-range training to suppress accumulation error and make the codec perform better on a long group of pictures (GOP). Overall, existing neural video coding methods improve compression efficiency from various perspectives, such as inter- or bi-directional prediction [30, 15, 14, 16, 39, 40], residual/context compression [13], and entropy models [29, 13]. Many NVC methods, such as ECVC-series [38] and DCVC-series [14, 2, 13, 15, 39], demonstrate excellent compression performance.
@@ II-B1 Analyze-then-Compress Paradigm
Figure 2: Overview of the proposed SEC-VCM framework. (1) The basic compression loop contains motion and contextual coding processes. “Q” denotes the quantization operation. Notably, some modules in this step are not shown above for simplicity, since this step is directly adopted from the neural video codec DCVC-HEM [29]. (2) The semantic decoding process is information-aligned with the encoding process of the pretrained visual backbone by the proposed BiEC mechanism from an entropy perspective. This helps the semantic decoder $\mathcal{D}_{\textit{sem}}$ purely handle information beneficial for MVS. (3) The pixel-level prior knowledge $f_{\textit{pixel}}$ is injected into the perception-oriented decoding process to fulfill frame content, suppress artifacts, and enhance final reconstruction quality.
@@ III-A Motivation and Overview
Existing neural video codecs typically adopt VAE-like networks [50] to compress the input video $x$ into compact latent representations and reconstruct the frame $\hat{x}$ for the machine vision system (MVS). Current methods optimize the codec for specific or multiple machine vision tasks, either in a supervised or unsupervised manner [4, 43, 5, 6]. Also, some methods maintain semantic consistency by minimizing the perceptual distance between the original video $x$ and compressed video $\hat{x}$ measured by visual backbone (VB) or visual foundation model (VFM) [8, 7]. Their paradigm is formulated as follows:
@@
where $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$ denote the encoder and decoder of the video codec, $\lfloor\cdot\rceil$” denotes the quantization operation. $\mathcal{L}$ denotes the total loss, and $R$ denotes the bitrate. $\mathcal{E}_{\textit{visual}}$ denotes the encoder of VB or VFM. $D(\cdot,\cdot)$ denotes pixel-domain or feature-domain distortion (e.g., MSE).
Although VB and VFM have been introduced to help codecs preserve semantics, their integration with video coding remains insufficiently explored. They excel at extracting rich and general feature representations for multiple machine vision tasks [21, 22], yet such features are not naturally suited for entropy coding or direct compression. To better leverage their general capacity, we align the video codec with VB/VFM from an entropy perspective, encouraging it to selectively preserve features beneficial for MVS while squeezing irrelevant information. The alignment objective is formulated as:
@@
where $\mathcal{D}_{\textit{sem}}(\cdot)$ denotes the semantic decoder and $\hat{y}$ is the transmitted bitstream (entropy-coded latent representation). Additionally, completing the compression loop requires low-level details to support inter-prediction [4]. Therefore, we propose that pixel-oriented reconstruction and machine-oriented reconstruction are conducted separately, with the former providing pixel-level priors for the latter.
Figure 3: Structures of proposed modules. (a) Structure of the information-alignment-based semantic decoder and BiEC mechanism. DEDE denotes the distribution estimation module, and $\uparrow\times 2$” denotes two times up-sampling. During training, the intermediate features for the decoding $f_{\textit{MVS}}$ are made hierarchically symmetric with the features $f_{\textit{VB}}$ from the visual backbone. (b) Structure of the SPDF module. Machine vision system (MVS)-oriented semantic feature $f_{t}^{\textit{MVS}}$ and pixel-related feature $f_{t}^{\textit{pixel}}$ are selectively combined based on the factor $\alpha$.
To achieve the above ideas, we propose the Symmetric Entropy-Constrained Video Coding framework for Machine (SEC-VCM), as shown in Fig. 2. Specifically, it consists of three main steps: inter-prediction and contextual coding for the basic compression loop, machine-oriented semantic decoding with the Bi-directional Entropy-Constraint (BiEC) mechanism, and semantic-pixel gated fusion. (1) In the first step, we inherit the basic compression loop from DCVC-HEM [29], where motion and contextual information are compressed, quantized, entropy-coded, and then reconstructed. The compensated features $c_{1}$, $c_{2}$, and $c_{3}$ serve as motion priors for both contextual coding and our semantic decoding. The motion latent $\hat{y}_{t}^{mv}$ and contextual latent $\hat{y}_{t}^{ctx}$ are then transmitted in the bitstream. (2) In the second step, the contextual latent representation $\hat{y}t^{ctx}$ is employed for semantic decoding, where the intermediate feature $f_{\textit{MVS}}$ for MVS-oriented reconstruction is symmetrically aligned with $f_{\textit{VB}}$ extracted by the visual backbone $\mathcal{E}_{\textit{visual}}$. Specifically, in the BiEC mechanism, the symmetry is enforced at three scales, and the alignment is achieved by constraining conditional entropy. (3) In the third step, since $f_{\textit{MVS}}$ mainly contains the semantic information, additional detailed pixel-level priors $f_{\textit{pixel}}$ are injected into the final reconstruction process to enrich frame details, suppress artifacts, and produce frames better suited for MVS. Notably, our framework is general across multiple downstream tasks, and we therefore remove the task loss $\mathcal{L}_{task(\hat{x})}$ that relies on specific labeled datasets. Our paradigm is illustrated as follows:
@@
where $\mathcal{D}_{\textit{sem}}(\cdot)$ denotes the proposed information-alignment-based semantic decoder, $f_{\textit{VB}}$ denotes the feature extracted by the visual backbone, $f_{\textit{MVS}}$ denotes the feature in the semantic decoder, and $f_{\textit{pixel}}$ denotes the feature for low-level reconstruction. $H(f_{\textit{MVS}}|f_{\textit{VB}})$ and $H(f_{\textit{VB}}|f_{\textit{MVS}})$ denote the conditional entropy terms used as loss items in the proposed BiEC mechanism.
@@ III-B Semantic Decoding with Symmetric Entropy-Constraint
To leverage the visual backbone’s strong capability of extracting machine-vision-oriented feature representations, the video decoding process is hierarchically and symmetrically aligned with the encoding process of VB during training. It enables the decoder to inherit the prior knowledge learned by the backbone that has experienced large-scale pretraining, focusing purely on features relevant to machine vision tasks. To achieve this idea, an information-aligned semantic decoder and a BiEC mechanism are designed, as shown in Fig. 3(a).
@@ III-B1 Information-Alignment-based Semantic Decoder
The semantic decoder comprises four bottleneck residual blocks and three up-sampling layers. It progressively recovers the quantized and compact latent representation $\hat{y}_{ctx}$ (at $1/16$ resolution) back to the original spatial resolution, producing machine-vision-oriented features $f_{\textit{MVS}}$. The inter-predicted features $c_{2}$ ($1/2$ size) and $c_{3}$ ($1/4$ size) are used as conditions for frame reconstruction, referring to [28, 29]. Notably, intermediate features $f_{\textit{MVS}}$ within the semantic decoder are aligned with the intermediate results $f_{\textit{VB}}$ from the visual backbone at three distinct scales. These three hierarchical alignments make the decoder symmetric to the pretrained guidance backbone.

### v2

@@ Abstract
As video transmission increasingly serves machine vision systems (MVS) instead of human vision systems (HVS), video coding for machines (VCM) has become a critical research topic. Existing VCM methods often bind codecs to specific downstream models, requiring retraining or supervised data, thus limiting generalization in multi-task scenarios. Recently, unified VCM frameworks have employed visual backbones (VB) and visual foundation models (VFM) to support multiple video understanding tasks with a single codec. They mainly utilize VB/VFM to maintain semantic consistency or suppress non-semantic information, but seldom explore how to directly link video coding with understanding under VB/VFM guidance. Hence, we propose a Symmetric Entropy-Constrained Video Coding framework for Machines (SEC-VCM). It establishes a symmetric alignment between the video codec and VB, allowing the codec to leverage VB’s representation capabilities to preserve semantics and discard MVS-irrelevant information. Specifically, a bi-directional entropy-constraint (BiEC) mechanism ensures symmetry between the process of video decoding and VB encoding by suppressing conditional entropy. This helps the codec to explicitly handle semantic information beneficial to MVS while squeezing useless information. Furthermore, a semantic-pixel dual-path fusion (SPDF) module injects pixel-level priors into the final reconstruction. Through semantic-pixel fusion, it suppresses artifacts harmful to MVS and improves machine-oriented reconstruction quality. Experimental results show our framework achieves state-of-the-art (SOTA) in rate-task performance, with significant bitrate savings over VTM on video instance segmentation (37.4%), video object segmentation (29.8%), object detection (46.2%), and multiple object tracking (44.9%). We will release our code soon.
@@ I Introduction
Figure 1: Overview of different methods. $I(\cdot)$ denotes mutual information, $H(\cdot)$ denotes Shannon entropy, and $\Uparrow$” denotes that the optimization goal is to maximize the value. $\mathcal{L}_{a}$ and $\mathcal{L}_{b}$ denote downstream task losses, $\mathcal{L}_{\textit{cons}}$ denotes loss for maintaining feature consistency, and $\mathcal{E}_{\textit{visual}}$ denotes the encoder of visual backbone (VB) or visual foundation model (VFM). (a) Standard or neural video codecs [1, 2] are optimized for HVS, which makes the decoded videos $\hat{x}$ suboptimal for machine vision. (b) Each neural video/feature codec [3, 4, 5] is optimized for a specific machine vision task by downstream task loss. (c) The unified codec [6, 7, 8, 9] is optimized with the assistance of the VB or VFM to reconstruct videos suitable for multiple downstream tasks. (d) Our unified compression framework aligns the video coding process with the video understanding process from the entropy perspective and squeezes machine-vision-irrelevant information during the process of video reconstruction.
Digital video plays a crucial role in our daily lives, comprising a significant portion of data storage. To efficiently store and transmit this massive amount of data, video coding techniques have been developed, primarily designed to meet the requirements of the human visual system (HVS). For videos aimed at HVS, such as movies and short clips, it is essential to preserve visual details perceptible to humans during compression. Meanwhile, videos collected for machine vision tasks, such as surveillance analysis [10] and facial recognition [11], typically require only part of the video information, rather than the complete visual details [12, 13]. Currently, neural video coding frameworks have evolved significantly and offer excellent compression performance [14, 15, 16, 17, 18, 19]. However, these methods are optimized primarily for pixel-domain distortion or HVS-related metrics and do not fully satisfy the requirements of machine vision systems (MVS) in terms of video understanding (e.g., detection, segmentation, and object tracking) [20], as shown in Fig. 1(a). Therefore, it is essential to develop video coding for machines (VCM) methods and bridge the gap between video coding and video understanding.
To reconstruct machine-vision-friendly videos, most existing methods [21, 22, 4] bridge the upstream codec with specific downstream task models and jointly optimize with the loss functions of different tasks, as shown in Fig. 1(b). However, this requires retraining the entire upstream codec or downstream models for each new task, resulting in significant computational and time costs. Meanwhile, some parameter-efficient fine-tuning (PEFT) strategies are introduced to adapt codecs to new downstream tasks with lower cost [3]. Nonetheless, most of them still rely on the labeled data and supervised training process for specific downstream tasks, limiting their generalization in multi-task scenarios.
@@
To address the above issues, we propose a Symmetric Entropy-Constrained Video Coding framework for Machine (SEC-VCM). It supports multiple downstream tasks with a single unsupervised training process and further eliminates machine-vision-irrelevant information by leveraging existing visual backbones, as shown in Fig. 1(d). Its key principle is to bridge the gap between the video coding and understanding processes from an entropy perspective, explicitly guiding the video codec to handle only information beneficial to video understanding while discarding irrelevant information. Specifically, an information-alignment-based semantic decoder with a bi-directional entropy-constraint (BiEC) mechanism is proposed to make the video decoding process of the codec symmetric and consistent with the encoding process of the visual backbone. In addition, a semantic-pixel dual-path fusion (SPDF) module is designed to incorporate HVS-related priors into the machine-oriented reconstruction process. It helps reduce artifacts harmful to MVS and fulfill the frame content through a gated fusion mechanism. The experiments on four machine vision tasks demonstrate that the proposed SEC-VCM achieves state-of-the-art (SOTA) rate-task performance. Compared to VTM, our method achieves significant bitrate reductions, with savings of 37.4% for video instance segmentation, 29.8% for video object segmentation, 46.2% for object detection, and 44.9% for multiple object tracking.
@@ II-A Neural Video Coding
Most existing neural video coding methods [29, 30, 31, 16, 32, 17, 18, 33] adopt a two-step pipeline including inter-prediction and residual or context compression. Lu et al. [29] proposed the first neural video codec DVC, which uses the optical flow [34] for inter-prediction. Hu et al. [35] proposed to convert the processes of motion estimation (ME), motion compensation (MC), and residual coding from the pixel domain to the feature domain. Li et al. [30] improved the concept of residual coding and proposed the conditional coding paradigm for better compression performance, termed DCVC. Sheng et al. [36] upgraded the DCVC and proposed DCVC-TCM with a multi-scale conditional coding to better remove temporal redundancy. Furthermore, many methods [17, 15, 19] introduce long-range training to suppress accumulation error and make the codec perform better on a long group of pictures (GOP). Overall, existing neural video coding methods improve compression efficiency from various perspectives, such as inter- or bi-directional prediction [32, 17, 16, 18, 37], residual/context compression [14], and entropy models [31, 14]. Many NVC methods, such as ECVC-series [19] and DCVC-series [16, 15, 14], demonstrate excellent compression performance.
@@ II-B1 Analyze-then-Compress Paradigm
Figure 2: Overview of the proposed SEC-VCM framework. (1) The basic compression loop contains motion and contextual coding processes. “Q” denotes the quantization operation. Notably, some modules in this step are not shown above for simplicity, since this step is directly adopted from the neural video codec DCVC-HEM [31]. (2) The semantic decoding process is information-aligned with the encoding process of the pretrained visual backbone by the proposed BiEC mechanism from an entropy perspective. This helps the semantic decoder $\mathcal{D}_{\textit{sem}}$ purely handle information beneficial to MVS. (3) The pixel-level prior knowledge $f_{\textit{pixel}}$ is injected into the perception-oriented decoding process to fulfill frame content, suppress artifacts, and enhance final reconstruction quality.
@@ III-A Motivation and Overview
Existing neural video codecs typically adopt VAE-like networks [46] to compress the input video $x$ into compact latent representations and reconstruct the frame $\hat{x}$ for the machine vision system (MVS). Current methods optimize the codec for specific or multiple machine vision tasks, either in a supervised or unsupervised manner [4, 5, 6, 7]. Also, some methods maintain semantic consistency by minimizing the perceptual distance between the original video $x$ and the compressed video $\hat{x}$ measured by visual backbone (VB) or visual foundation model (VFM) [9, 8]. Their paradigm is formulated as follows:
@@
where $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$ denote the encoder and decoder of the video codec, $\lfloor\cdot\rceil$” denotes the quantization operation [46]. $\mathcal{L}$ denotes the total loss, $R$ denotes the bitrate, $\mathcal{E}_{\textit{visual}}$ denotes the encoder of VB or VFM, $D(\cdot,\cdot)$ denotes pixel-domain or feature-domain distortion (e.g., MSE), and $\mathcal{L}_{task(\hat{x})}$ denotes the loss function of the downstream task.
Although VB and VFM have been introduced to help codecs preserve semantics, their integration with video coding remains unexplored. They excel at extracting rich and general feature representations for multiple machine vision tasks [24, 25], yet such features are not naturally suited for entropy coding or direct compression. To better leverage their general capacity, we propose to align the video codec with VB/VFM from an entropy perspective, encouraging it to selectively preserve features beneficial to MVS while squeezing irrelevant information. The alignment objective is formulated as:
@@
where $\mathcal{D}_{\textit{sem}}(\cdot)$ denotes the semantic decoder and $\hat{y}$ denotes the transmitted bitstream (entropy-coded latent representation). Additionally, completing the compression loop requires low-level details to support inter-prediction [4]. Therefore, we propose that pixel-oriented reconstruction and machine-oriented reconstruction are conducted separately, with the former providing pixel-level priors for the latter.
Figure 3: (a) Structure of the proposed information-alignment-based semantic decoder and BiEC mechanism. DEDE denotes the distribution estimation module, and $\uparrow\times 2$” denotes two times up-sampling. During training, the intermediate features for the decoding $f_{\textit{MVS}}$ are made hierarchically symmetric with the features $f_{\textit{VB}}$ from the visual backbone. (b) Structure of the proposed SPDF module. Machine vision system (MVS)-oriented semantic feature $f_{t}^{\textit{MVS}}$ and pixel-related feature $f_{t}^{\textit{pixel}}$ are selectively combined based on the factor $\alpha$.
To achieve the above ideas, we propose the Symmetric Entropy-Constrained Video Coding framework for Machine (SEC-VCM), as shown in Fig. 2. Specifically, it consists of three main steps: inter-prediction and contextual coding for the basic compression loop, machine-oriented semantic decoding with the Bi-directional Entropy-Constraint (BiEC) mechanism, and semantic-pixel gated fusion. (1) In the first step, we adopt the basic compression loop from DCVC-HEM [31], where motion and contextual information are compressed, quantized, entropy-coded, and then reconstructed. The compensated features $c_{1}$, $c_{2}$, and $c_{3}$ serve as motion priors for both contextual coding and our semantic decoding. The motion latent $\hat{y}_{t}^{mv}$ and contextual latent $\hat{y}_{t}^{ctx}$ are then transmitted in the bitstream. (2) In the second step, the contextual latent representation $\hat{y}t^{ctx}$ is employed for semantic decoding, where the intermediate feature $f_{\textit{MVS}}$ for MVS-oriented reconstruction is symmetrically aligned with $f_{\textit{VB}}$ extracted by the visual backbone $\mathcal{E}_{\textit{visual}}$. Specifically, in the BiEC mechanism, the symmetry is enforced at three scales, and the alignment is achieved by constraining conditional entropy. (3) In the third step, since $f_{\textit{MVS}}$ mainly contains the semantic information, additional detailed pixel-level priors $f_{\textit{pixel}}$ are injected into the final reconstruction process to enrich frame details, suppress artifacts, and produce frames better suited to MVS. Notably, our framework is general across multiple downstream tasks; therefore, we remove the task loss $\mathcal{L}_{task(\hat{x})}$ that relies on specific labeled datasets. Our paradigm is illustrated as follows:
@@
where $\mathcal{D}_{\textit{sem}}(\cdot)$ denotes the proposed information-alignment-based semantic decoder, $f_{\textit{VB}}$ denotes the feature extracted by the visual backbone, $f_{\textit{MVS}}$ denotes the feature in the semantic decoder, and $f_{\textit{pixel}}$ denotes the feature used for low-level reconstruction. $H(f_{\textit{MVS}}|f_{\textit{VB}})$ and $H(f_{\textit{VB}}|f_{\textit{MVS}})$ denote the conditional entropy terms used as loss items in the proposed BiEC mechanism.
@@ III-B Semantic Decoding with Symmetric Entropy-Constraint
To leverage the visual backbone’s strong capability for extracting machine-vision-oriented feature representations, the video decoding process is hierarchically and symmetrically aligned with the encoding process of VB during training. It enables the decoder to inherit the prior knowledge learned by the backbone that has experienced large-scale pretraining, focusing purely on features relevant to machine vision tasks. To achieve this idea, an information-aligned semantic decoder and a BiEC mechanism are designed, as shown in Fig. 3(a).
@@ III-B1 Information-Alignment-based Semantic Decoder
The semantic decoder comprises four bottleneck residual blocks and three up-sampling layers. It progressively recovers the quantized and compact latent representation $\hat{y}_{ctx}$ (at $1/16$ resolution) back to the original spatial resolution, producing machine-vision-oriented features $f_{\textit{MVS}}$. The inter-predicted features $c_{2}$ (at $1/2$ resolution) and $c_{3}$ (at $1/4$ resolution) are used as conditions for frame reconstruction, referring to [30, 31]. Notably, intermediate features $f_{\textit{MVS}}$ within the semantic decoder are aligned with the intermediate results $f_{\textit{VB}}$ from the visual backbone at three distinct scales. These three hierarchical alignments make the decoder symmetric to the pretrained guidance backbone.